A boy is now one-third as old as his father
Twelve years hence he will be half as old as
his father. Then the present age of the boy is

Found 2 solutions by lwsshak3, clettol:
Answer by lwsshak3(11628)

A boy is now one-third as old as his father
Twelve years hence he will be half as old as
his father. Then the present age of the boy is
***
let x=age of boy
3x=age of father
..
12 yrs hence:
x+12=age of boy
3x+12=age of father
3x+12=2(x+12)
3x+12=2x+24
x=12
present age of the boy is 12

Answer by clettol(1)

3x =age of the father
12 = age of the boy
x+12=age of the boy
3x+12=2(X+12)
3X+12=2X+24

X=12